OpenStudy (anonymous):

What is 8/10 minus 2/3

OpenStudy (cggurumanjunath):

8/10 =(8/2) / (10/2) =4/5

OpenStudy (cggurumanjunath):

4/5-2/3=(4*3-5*2)/(5*3)

OpenStudy (cggurumanjunath):

(12-10)/15=2/15